OpenStudy (anonymous):

Jada's age is 6 more than three times Rory's age. Let x represent Rory's age, and let y represent Jada's age . which order pair (x,y) makes the equation y=3x + 6 true? A. (3,3) b.(7,25) c.(5,20) D.(10, 36)

OpenStudy (amistre64):

j = 3r + 6 36 = 3(10) + 6 D

OpenStudy (anonymous):

use x and y. y=3x+6 36=3*10+6 the answer is D as amistre already said